# Don’t be Fooled: The Misinformation Effect of Explanations in Human-AI Collaboration

### v1

@@ Appendix A Appendix
The table shows the prompts used to generated the output of the LLM. Prompt for correct explanations Hello, you have to provide support to a human in classifying the architectural style of buildings in images. Overall, there are three different styles to choose from: Art Deco, Art Nouveau and Georgian Architecture. For an image, please also provide within two to three sentences an explanation for your decision. In the following is the image: Prompt for incorrect explanations Hello, you have to classify the architectural style of buildings in images. Overall, there are three different styles to choose from: Art Deco, Art Nouveau, and Georgian Architecture. Please provide the classification of an image. Additionally, we want to investigate how false explanations can mislead users. Thus, also explain your decision within 2-3 sentences. However, make sure that the explanation is incorrect, thus it does not correspond to your classification. Can you please provide a misleading explanation by using actual elements of the image? Here is the image:
@@
The table shows the results of the regression model with the main task performance as the dependent variable. In the columns, the coefficient and the standard error are presented. As significance levels, .1, .05, and 0.01 are defined. Dependent variable Performance Coeff SE Intercept .676*** .083 condition_AI [True] .259*** .039 condition_correctexplanation [True] .311*** .040 condition_incorrectexplanation [True] .250*** .038 cognitive_load -.023 .087 AI trust -.237* .134 AI knowledge -.029 .061 AI attitude .092 .096 .367 Adj. .338 Log-Likelihood 59.892 F-statistic 12.580***
@@
The table shows the results of the regression model with the post-test performance as the dependent variable. In the columns, the coefficient and the standard error are presented. As significance levels, .1, .05, and 0.01 are defined. Dependent variable Performance Coeff SE Intercept .618*** .112 condition_AI [True] .125** .053 condition_correctexplanation [True] .198*** .054 condition_incorrectexplanation [True] .117** .052 cognitive_load -.197* .118 AI trust -.136 ,180 AI knowledge .006 .082 AI attitude .033 .130 .127 Adj. .087 Log-Likelihood 11.914 F-statistic 3.153***

### v2

@@ Appendix A Appendix
The table shows the prompts used to generated the output of the LLM. Prompt for correct explanations Hello, you have to provide support to a human in classifying the architectural style of buildings in images. Overall, there are three different styles to choose from: Art Deco, Art Nouveau and Georgian Architecture. For an image, please also provide within two to three sentences an explanation for your decision. In the following is the image: Prompt for incorrect explanations Hello, you have to classify the architectural style of buildings in images. Overall, there are three different styles to choose from: Art Deco, Art Nouveau, and Georgian Architecture. Please provide the classification of an image. Additionally, we want to investigate how false explanations can mislead users. Thus, also explain your decision within 2-3 sentences. However, make sure that the explanation is incorrect, thus it does not correspond to your classification. Can you please provide a misleading explanation by using actual elements of the image? Here is the image: Prompt for reasoning scores Please Provide an assessment of the correctness of the human explanation, and give the response back with only a correctness score for each human explanation. Here is a human explanation and ground truth explanation regarding three different architectural styles: Art Nouveau, Art Deco, and Georgian Architecture. The human explanation is divided into three parts in the order of Art Nouveau, Art Deco, and Georgian Architecture. Human Explanation: Ground Truth: . All Ground truths can be seen as high-score explanations. Correctness refers to the truthfulness/faithfulness of the explanation. Decide on a scale of 0-5 for scoring explanations: 0: Completely incorrect or irrelevant. 1: Major errors or omissions; very limited correct information. 2: Some correct information, but with significant errors or incomplete coverage. 3: Generally correct but with minor inaccuracies or omissions. 4: Mostly accurate and complete, with minor errors. 5: Completely correct and comprehensive.
@@
The table shows the results of the regression model with the main task performance as the dependent variable. In the columns, the coefficient and the standard error are presented. As significance levels, .1, .05, and 0.01 are defined. Dependent variable Performance Coeff SE Intercept .676*** .083 condition_AI [True] .259*** .039 condition_correctexplanation [True] .311*** .040 condition_incorrectexplanation [True] .250*** .038 cognitive_load -.023 .087 AI trust -.237* .134 AI knowledge -.029 .061 AI usefulness .092 .096 .367 Adj. .338 Log-Likelihood 59.892 F-statistic 12.580***
@@
The table shows the results of the regression model with the post-test performance as the dependent variable. In the columns, the coefficient and the standard error are presented. As significance levels, .1, .05, and 0.01 are defined. Dependent variable Performance Coeff SE Intercept .618*** .112 condition_AI [True] .125** .053 condition_correctexplanation [True] .198*** .054 condition_incorrectexplanation [True] .117** .052 cognitive_load -.197* .118 AI trust -.136 ,180 AI knowledge .006 .082 AI usefulness .033 .130 .127 Adj. .087 Log-Likelihood 11.914 F-statistic 3.153***
@@
Randomly selected answers to the open-ended questions. Participant Answers Reasoning score Art Nouveau Art Deco Georgian P7 Art Nouveau draws inspiration for nature in lines and form while still maintaining modernity and simplicity in building materials. The flow of the lines complement rigid forms and lines, and assymetry is celebrated. Art Deco is all about geometry, curve, and sleek modernism (as interpreted in the early 20th century). The lines are smooth and fluid, yet with some repetition that is very pleasing. Symmetry is celebrated and enhanced with that same repetition, and you can get a sense of strong and soaring feelings in some of the buildings and compact sleekness in others. Georgian Architecture is a late 17th-early 18th century style that celebrates symmetry, repetition, and plain lines sometimes enhanced with masonry features also very strong and simple in their form. Brick, stone, and stucco are very much in evidence. The entrance is a dominant feature, often with a fan window above the door and column or column-like surrounds with decorative capitals. 3.77 P34 I would define architectural style Art Nouveau as the form of architecture that incorporate natural and historic features in environment in its design. Architecturral style Art Deco is the form of architecture that incorporates a lot of angular and triangular designs and are also very colourful. Georgian Architecture is a form of architecture that incorporates a lot of windows and doors in its design. These door and windows are fixed at a distance from the wall. 2.0 P74 flowing, organic shapes inspiMaroon by nature, curving lines, assymetry, and the use of materials like iron, glass, and ceramics. The design often includes decorative, whimsical elements. angular, upward-pointing lines, geometric shapes, and a combination of opulent materials like stucco, terra-cotta, and metals. The style often features bold colors and sharp contrasts. symmetrical facades with a balanced arrangement of windows and doors. Features include fanlights above doors and materials like brick, stone, and wood. Georgian buildings are known for their classical proportions and solid, enduring construction. 4.33 P135 Art Nouveau seems the most modern of the 3 and can take various somewhat different looks. Art Deco seems very old style. With concrete buildings that are somewhat tall Georgian Architecture has a specific look that seems to have bricks very often and buildings that do not look modern at all. 1.66
The table shows the results of the regression model with the reasoning results as the dependent variable. In the columns, the coefficient and the standard error are presented. As significance levels, .1, .05, and 0.01 are defined. Dependent variable Reasoning scores Coeff SE Intercept 2.611*** .349 condition_AI [True] .171 .165 condition_correctexplanation [True] .058 .169 condition_incorrectexplanation [True] -.144 .161 cognitive_load -.109 .367 AI trust -.712 .561 AI knowledge -.192 .257 AI usefulness .744* .404 .066 Adj. .023 Log-Likelihood -169.82 F-statistic 1.540
1
Note: * p ¡ .1; ** p ¡ .05; *** p ¡ .01